OpenStudy (anonymous):

If the polynomial \(kx^2 - 2xy - 3y^2 + 3x - 5y +2 \) can be factorized into product of two linear factors, what is k?

OpenStudy (anonymous):

(x-3y+1)(x+y+2) where k=1

OpenStudy (anonymous):

Looking at the terms, I decided that they had to be two factors each with (ax+by+c). I started by thinking that the trailing term of 2 must come from 1 and 2. I looked at the -3y^2 term and tried to come up with different ways to get it if the result gave me -5y down the road. Same with -2xy. I used trial and error with 3 different sets of numbers, each time tweaking a minus sign or two coefficients until I say how the pattern was coming out. I wish I had a better analytical method.

OpenStudy (anonymous):

Trial and error :( First, factorize -3y^2 - 5y + 2 -3y^2 - 5y + 2 = (y+2)(-3y+1) Let kx^2 - 2xy -3y^2 + 3x - 5y +2 = (lx + y+2)(mx - 3y + 1) = lm x^2 + (m-3l)xy - 3y^2 + (l+2m)x -5y + 2 By equating the coefficients, lm = k m - 3l = -2 2m + l = 3 => l = 1, m = 1 , k = lm = 1 Any better method?

OpenStudy (anonymous):

I like that much better. I was thinking that with my method, if k was something other than 1 I would be in trouble. Your method seems to cut through that.

OpenStudy (anonymous):

one can let\[kx^2−2xy−3y^2+3x−5y+2=(kx+ay+1)(x+by+2)\]and solve for a,b,k...without trial and error but this is not easier, is it?

OpenStudy (anonymous):

allow me to do some effort

OpenStudy (anonymous):

Another way I've tried, but I don't know what I am actually doing: kx^2 - 2xy - 3y^2 + 3x - 5y + 2 = 0 -3y^2 - (5+2x)y + kx^2 + 3x + 2 = 0 Δ = [-(5+2x)]^2 - 4 (-3)(kx^2 + 3x + 2) = 25 + 20x + 4x^2 + 12kx^2 + 36x + 24 = (4+12k)x^2 + 56x + 49 Since it can be factored, it should be in (ax+b)^2, and b is 7, a = 56/(7x2) = 4 So, \(\sqrt{4+12k} = 4\) 4 + 12k = 16 k = 1

OpenStudy (anonymous):

after simplifying\[kx^2−2xy−3y^2+3x−5y+2\\=kx^2+(a+bk)xy+aby^2+(2k+1)x+(2a+b)y+2\]2k+1=3--->k=1 a+b=-2 2a+b=-5---->a=-3 , b=1

OpenStudy (anonymous):

Allow me to type something here. \[(A_1x + B_1y+C_1)(A_2x + B_2y+C_2)\]\[=(A_1x + B_1y)(A_2x + B_2y)+C_1(A_2x + B_2y)+C_2(A_1x + B_1y) +C_1C_2\] \((A_1x + B_1y)(A_2x + B_2y)\) contains \(x^2, xy, y^2\) terms; \(C_1(A_2x + B_2y)+C_2(A_1x + B_1y)\) contains x and y terms; \(C_1C_2\) is a constant. So, we can factorize the x^2, xy and y^2 terms first,

OpenStudy (anonymous):

By the way, does my second method make sense?

OpenStudy (anonymous):

some trial and error used for last step but makes sense :)

OpenStudy (anonymous):

Trial and error? Where? Why?

OpenStudy (anonymous):

lol...wait..its accurate

OpenStudy (anonymous):

Wait. First method = first comment Second method = Δ one ''Something'' that I typed = background for the first method.

OpenStudy (anonymous):

Hmm, I guess I mixed that ''something'' up with the previous question. In this question, the term y^2, y and constant are factorized first as k, which is the coefficient of x^2, is the unknown.

OpenStudy (anonymous):

Look like I don't really understand the first method :(

OpenStudy (anonymous):

I used the first method again, but with a few more variables (allowing k to be a product other than (k)(1). It was a total mess with a couple mistakes that I had to find, but I got the original answer along with (2x-6y+2)(0.5x+0.5y+1) where k=1. No trial and error. Thank you for teaching me a new skill.

OpenStudy (anonymous):

You're welcome! Actually, I just learnt it today :P